OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57 21 -21 9 -9

OpenStudy (anonymous):

If you double the second equation, you obtain: 8x + 2y = 114. Note the first equation, too: 7x - 2y = 21 8x + 2y = 114. If this is not suggestive of anything, look at the y values. What I'm trying to say is to add the two equations: 15 x = 135. x = 9